# Limited Memory LRSGA Optimizer to competitive optimization

### v1

@@ 4.1 Experiment on Mnist
The first experiments is conducted on the MNIST dataset, a classic benchmark in machine learning and computer vision, consisting of 70,000 grayscale images of handwritten digits from 0 to 9. Each image is $28\times 28$ pixels and centered within the frame to standardize positioning. The data set is divided into 60,000 training samples and 10,000 test samples, and its simplicity, small size, and well-defined labels have made it a fundamental starting point for testing and comparing classification algorithms. We used only the training split to train the GAN.
@@ 4.1.1 Results
Figure 1(a) presents the Fréchet Inception Distance (FID) scores over training epochs on the MNIST dataset. It compares our proposed optimizer LM-LRSGA, and its exponential moving average variant (LRSGAEma) against the widely used Adam optimizer. Both LM-LM-LRSGAand LM-LRSGAEma achieve substantially lower FID scores throughout training than Adam, demonstrating superior sample quality. While Adam begins with a significantly higher FID score (above 70) and gradually decreases to around 19 after 100 epochs, LM-LRSGAand LM-LRSGAEma start with lower scores ($\sim 45$) and converge much faster. Notably, LM-LRSGAEma consistently outperforms LRSGA, yielding smoother convergence and the lowest final FID score of approximately 13. These results highlight the effectiveness of our optimizer in stabilizing generative model training and producing higher-quality samples compared to the conventional Adam baseline.
Figure 1(b) shows the generator and discriminator loss trajectories on the MNIST dataset when trained with LRSGA, LM-LRSGAEma, and Adam. The results indicate that LM-LRSGAand LM-LRSGAEma lead to more stable and balanced adversarial training than Adam. With Adam, the generator loss remains consistently lower ($\sim 1.3$) while the discriminator stabilizes around $\sim 0.8$, reflecting a tendency toward imbalance and potential mode collapse. In contrast, LM-LRSGAand LM-LRSGAEma exhibit more synchronized dynamics: the generator loss converges to $\sim 1.4$, while the discriminator loss steadily decreases toward $\sim 0.5$, suggesting healthier competition between the two networks. Notably, LM-LRSGAEma achieves the smoothest convergence, stabilizing both losses earlier and exhibiting minimal oscillation. These findings align with the FID results, demonstrating that our proposed optimizer fosters more stable GAN training and enhances generative performance relative to the Adam baseline.
@@
Figure 1: Comparison of LM-LRSGAEma and Adam optimizers on MNIST datasets.
@@ 4.1.2 Sensitivity analysis
In this section, we analyze the role of the main parameters of the LRMSGAEma algorithm. We studied how the model ability generates realistic images when we change the value of a parameter.
Figure 2(a) presents the hyperparameter sensitivity analysis $\eta$ on the minimum Fréchet Inception Distance (FID) score achieved during the experiments. The bar chart shows results for $\eta$ values ranging from 0.05 to 0.30. Performance varies notably across settings, with the highest (worst) FID score observed at $\eta=0.05$, indicating degraded generative quality in this regime. The best results are obtained for $\eta$ values between 0.20 and 0.25, yielding minimum FID scores around 14, suggesting an optimal trade-off between learning stability and image fidelity. Larger $\eta$ values beyond 0.25 slightly worsen the FID, though still outperforming the smallest $\eta$ tested. This pattern highlights the importance of tuning $\eta$ for maximizing GAN performance.
Figure 2(b) reports the hyperparameter sensitivity analysis for the parameter $\tau$. The results are shown for $\tau$ values from 0 to 0.2. The 0 value is an experiment in which the component is ignored. Across this range, the minimum FID scores vary only slightly, indicating that the model performance is relatively insensitive to $\tau$ within the tested limits. The best generative quality is observed at $\tau=0.0002$, both achieving FID scores just above 16, while the highest scores (worst performance) occur around $\tau=0.002$. In general, the flat trend suggests that $\tau$ has a limited effect on GAN performance in this setup, allowing flexibility in its selection without a major impact on output fidelity.
Figures 2(c) and 2(d) show the results of the experiments to evaluate the effect of $\epsilon$ and max_history parameters. The charts show only negligible differences between the settings. For the $\epsilon$ parameter, we expect its impact to be minimal, as it is effectively offset by the residuals observed between consecutive optimization iterations. As for the max_history parameter, we consider a value of 10 sufficient for this dataset to accurately estimate the correction to be applied and the resource usage required to execute the experiment.
@@ 4.2 Experiment on Fashion Mnist
In this section we analyze the results about the Fashion-MNIST dataset, it is a widely used benchmark for image classification tasks, designed as a more challenging drop-in replacement for the original MNIST handwritten digits. It consists of 70,000 grayscale images of size $28\times 28$ pixels, split into 60,000 training samples and 10,000 test samples. We used only the training split to train the GAN models. Each image depicts a single fashion item from one of 10 categories, including T-shirts, trousers, pullovers, dresses, etc. The dataset compact size and diversity make it a standard testbed for evaluating machine learning algorithms in vision research.
Figure 3(a) shows the evolution of the FID scores over 100 training epochs for the proposed LM-LRSGAand LM-LRSGAEma optimizer and the Adam baseline optimizer on the Fashion dataset. The FID score decreases sharply for all methods during the initial epochs, dropping from above 90 to approximately 60–65 by epoch 20, with Adam slightly outperforming LM-LRSGAEma in this early phase. However, from epoch 30 onward, LM-LRSGAEma consistently achieves lower FID scores, and the performance gap increases as training progresses; while the LM-LRSGAachieves best performance across all epochs. In the final epoch, LM-LRSGAEma reaches a FID of around 30, compared to Adam’s $\sim 35$, and even surpassing LRSGA; indicating higher image fidelity. These results suggest that, while Adam converges rapidly at the start, LM-LRSGAEma delivers more stable improvements and better final generative quality.
Figure 3(b) shows the training losses of Adam and LM-LRSGAEma. The same behavior of Mnist dataset is observed in this experiment. The two losses converge to a stable value around the epoch $\sim 30$ for LM-LRSGAEma, while Adam converges to some epochs before. The distance between the two losses confirms the ability of LM-LRSGAEma to better deal than Adam.
@@
Figure 3: Comparison of LRSGA, LM-LRSGAEma, and Adam optimizers on Fashion datasets.

### v2

@@ 4.2.2 Transpose Two-loop recursion
5 On the local convergence of LM-LRSGA
In this section we prove a local linear convergence guarantee for the limited–memory variant introduced in Section 4. Specifically, for any fixed history length $\ell$, and for sufficiently small stepsizes $\eta$ and $\tau$, the LM-LRSGA iterates converge linearly to a stable Nash equilibrium, with constants depending on $\ell$.
Throughout the section we work on a convex open set $\Omega\subset\mathbb{R}^{m+n}$ and assume $f,g\in C^{3}(\Omega,\mathbb{R})$, so that the game gradient $F(w)=(\partial_{x}f(w),\partial_{y}g(w))^{\top}$ is of class $C^{2}$.
Recall that the LM-LRSGA iteration can be written as
$w_{k+1}=w_{k}-\eta\bigl(I-\tau\,\alpha_{k}^{(\ell)}\bigr)F(w_{k}),$ (20)
where $\eta>0$ and $\tau>0$ are fixed stepsizes and $\alpha_{k}^{(\ell)}$ is the limited–memory approximation of the antisymmetric component of the game Hessian
$\alpha_{k}^{(\ell)}:=\begin{pmatrix}0&\tfrac{1}{2}\bigl(M_{k}^{(\ell)}-(N_{k}^{(\ell)})^{\top}\bigr)\\[5.69054pt] \tfrac{1}{2}\bigl(N_{k}^{(\ell)}-(M_{k}^{(\ell)})^{\top}\bigr)&0\end{pmatrix}$, (21)
obtained from the truncated mixed–blocks $M_{k}^{(\ell)}$ and $N_{k}^{(\ell)}$ (see 18 and 19) with curvature pairs $(s_{k}^{x},s_{k}^{y},y_{k}^{f},y_{k}^{g})$ and fixed positive parameters $\varepsilon_{x},\varepsilon_{y}$.
We start by showing, in the following lemma, that the limited-memory mixed blocks $M_{k}^{(\ell)}$ and $N_{k}^{(\ell)}$ are uniformly bounded.
Lemma 5.1.
Assume that $\partial_{x}f$ and $\partial_{y}g$ are Lipschitz on $\Omega$ with constants $L_{x},L_{y}>0$. Then, for every fixed history length $\ell\geq 1$ there exist constants
$C_{M}^{x}(\ell):=(\ell+1)\bigl(L_{x}+|\varepsilon_{x}|\bigr),\qquad C_{M}^{y}(\ell):=(\ell+1)\bigl(L_{y}+|\varepsilon_{y}|\bigr),$
such that, for all $k$ with $s_{k}^{w}\neq 0$,
$\|M_{k}^{(\ell)}\|_{2}\leq C_{M}^{x}(\ell),\qquad\|N_{k}^{(\ell)}\|_{2}\leq C_{M}^{y}(\ell),$
where $s_{k}^{w}:=(s_{k}^{x},s_{k}^{y})$.
Proof.
Fix $k\geq 1$. By Lipschitz continuity of $\partial_{x}f$ on $\Omega$,
$\|\partial_{x}f(w_{k+1})-\partial_{x}f(w_{k})\|_{2}\leq L_{x}\|w_{k+1}-w_{k}\|_{2}=L_{x}\|s_{k}^{w}\|_{2}.$
Using the definition of $y_{k}^{f}$ and the inequality $\|s_{k}^{x}\|_{2}\leq\|s_{k}^{w}\|_{2}$ we obtain
$\|y_{k}^{f}\|_{2}\leq\|\partial_{x}f(w_{k+1})-\partial_{x}f(w_{k})\|_{2}+|\varepsilon_{x}|\,\|s_{k}^{x}\|_{2}\leq\bigl(L_{x}+|\varepsilon_{x}|\bigr)\,\|s_{k}^{w}\|_{2}.$
Recalling that $p_{k}:=1/\|s_{k}^{w}\|_{2}^{2}$ and considering that $\|s_{k}^{y}\|_{2}\leq\|s_{k}^{w}\|_{2}$, the rank–one term satisfies
$\|p_{k}y_{k}^{f}(s_{k}^{y})^{\top}\|_{2}\leq p_{k}\|y_{k}^{f}\|_{2}\,\|s_{k}^{y}\|_{2}\leq\frac{\bigl(L_{x}+|\varepsilon_{x}|\bigr)\,\|s_{k}^{w}\|_{2}\,\|s_{k}^{y}\|_{2}}{\|s_{k}^{w}\|_{2}^{2}}\leq L_{x}+|\varepsilon_{x}|.$
An identical argument applied to $\partial_{y}g$ and $y_{k}^{g}$ yields
$\|p_{k}y_{k}^{g}(s_{k}^{x})^{\top}\|_{2}\leq L_{y}+|\varepsilon_{y}|.$
The limited–memory formulas (18) and (19) express $M_{k}^{(\ell)}$ and $N_{k}^{(\ell)}$ as finite sums of at most $\ell+1$ such rank–one terms, multiplied on the right by products of the matrices
$\tilde{V}_{t}:=I-p_{t}\,s_{t}^{y}(s_{t}^{y})^{\top},\qquad V_{t}:=I-p_{t}\,s_{t}^{x}(s_{t}^{x})^{\top},$
for indices $t$ in the history window. Since $\|s_{t}^{x}\|_{2},\|s_{t}^{y}\|_{2}\leq\|s_{t}^{w}\|_{2}$ and $p_{t}=\|s_{t}^{w}\|_{2}^{-2}$, the eigenvalues of $\tilde{V}_{t}$ and $V_{t}$ lie in $[0,1]$, hence $\|\tilde{V}_{t}\|_{2}\leq 1$ and $\|V_{t}\|_{2}\leq 1$ for all $t$. The initializations $H_{0}^{k}$ (see Equation(17)) are also of the same form, and therefore satisfy the same bounds.
Then, by the triangle inequality we conclude that
$\|M_{k}^{(\ell)}\|_{2}\leq(\ell+1)\,\bigl(L_{x}+|\varepsilon_{x}|\bigr)=C_{M}^{x}(\ell),$
and analogously
$\|N_{k}^{(\ell)}\|_{2}\leq(\ell+1)\,\bigl(L_{y}+|\varepsilon_{y}|\bigr)=C_{M}^{y}(\ell),$
for every $k$ with $s_{k}^{w}\neq 0$. ∎
By denoting $\Delta_{k}^{(\ell)}:=\alpha_{k}^{(\ell)}-A^{*},$ the next corollary shows that $\Delta_{k}^{(\ell)}$ remains uniformly bounded along the iterations.
Corollary 5.2.
Under the assumptions of Lemma 5.1, let $w^{*}$ be a stable Nash equilibrium and denote by $A^{*}$ the antisymmetric part of the game Hessian at $w^{*}$. Then there exists a constant $\delta^{(\ell)}>0$ such that
$\|\alpha_{k}^{(\ell)}-A^{*}\|_{2}\leq\delta^{(\ell)}\qquad\text{for all }k.$
In particular, we may take
$C_{\alpha}(\ell):=\tfrac{1}{2}\bigl(C_{M}^{x}(\ell)+C_{M}^{y}(\ell)\bigr),\qquad\delta^{(\ell)}:=C_{\alpha}(\ell)+\|A^{*}\|_{2}.$
Proof.
As shown in Vater et al. (2025), for any rectangular matrix $C$ the block–skew matrix $B=\begin{pmatrix}0&C\\ -C^{\top}&0\end{pmatrix}$ satisfies $\|B\|_{2}=\|C\|_{2}$. Applying this to the definition (21) we obtain
$\|\alpha_{k}^{(\ell)}\|_{2}=\tfrac{1}{2}\bigl\|M_{k}^{(\ell)}-(N_{k}^{(\ell)})^{\top}\bigr\|_{2}\leq\tfrac{1}{2}\bigl(\|M_{k}^{(\ell)}\|_{2}+\|N_{k}^{(\ell)}\|_{2}\bigr)\leq C_{\alpha}(\ell)$
for all $k$, where $C_{\alpha}(\ell):=\tfrac{1}{2}\bigl(C_{M}^{x}(\ell)+C_{M}^{y}(\ell)\bigr)$. Finally,
$\|\Delta_{k}^{(\ell)}\|_{2}:=\|\alpha_{k}^{(\ell)}-A^{*}\|_{2}\leq\|\alpha_{k}^{(\ell)}\|_{2}+\|A^{*}\|_{2}\leq C_{\alpha}(\ell)+\|A^{*}\|_{2}=\delta^{(\ell)},$
which proves the claim. ∎
We finally combine Corollary 5.2 with the spectral condition of Section 3 to obtain a local convergence result for LM-LRSGA as claimed in the following theorem.
Letting $w^{*}\in\Omega$ be stable Nash equilibrium, for each $k$ and fixed history length $\ell$, we can define the $k$-frozen LM-LRSGA map
$T_{k,(\ell)}^{LM-LRSGA}(w):=w-\eta\,(I-\tau\alpha_{k}^{(\ell)})\,F(w),$
so that $w_{k+1}=T_{k,(\ell)}^{LM-LRSGA}(w_{k})$. Its Jacobian at the equilibrium $w^{*}$ is
$DT_{k,(\ell)}^{LM-LRSGA}(w^{*})=I-\eta G_{k}^{(\ell)},\qquad G_{k}^{(\ell)}:=(I-\tau\alpha_{k}^{(\ell)})\,H^{*}.$
In particular, the following theorem shows that, suitable choices of $\tau$ and $\eta$ yield local linear convergence of the full LM-LRSGA iteration.
Theorem 5.3.
Let $\Omega\subset\mathbb{R}^{m+n}$ be a convex open set and $f,g\in C^{3}(\Omega,\mathbb{R})$. Assume that $\partial_{x}f$ and $\partial_{y}g$ are Lipschitz on $\Omega$ with constants $L_{x},L_{y}>0$, and let $w^{*}\in\Omega$ be a stable Nash equilibrium. Then, for any fixed history length $\ell\geq 1$ there exist constants
$\bar{\tau}(\ell)>0,\qquad\bar{\eta}(\ell)>0,\qquad r(\ell)>0,$
such that the following holds. If $0<\tau<\bar{\tau}(\ell)$, $0<\eta<\bar{\eta}(\ell)$ and $w_{0}\in B_{r(\ell)}(w^{*})$, then:
  1. 1.
    the LM-LRSGA iterates $w_{k+1}=T_{k,(\ell)}^{LM-LRSGA}(w_{k})$ are well-defined and remain in $B_{r(\ell)}(w^{*})$ for all $k$;
  2. 2.
    the sequence $(w_{k})$ converges linearly to $w^{*}$, i.e. there exists a constant $q\in(0,1)$ such that
    $\|w_{k}-w^{*}\|_{2}\leq q^{k}\,\|w_{0}-w^{*}\|_{2}\qquad\text{for all }k\geq 0.$
Proof.
By Corollary 5.2, for each fixed $\ell$ there exists $\delta^{(\ell)}>0$ such that $\|\alpha_{k}^{(\ell)}-A^{*}\|_{2}\leq\delta^{(\ell)}$ for all $k$. Applying Corollary 3.3 with $\Delta_{k}^{(\ell)}:=\alpha_{k}^{(\ell)}-A^{*}$ we obtain a bound
$0<\tau<\bar{\tau}(\ell):=\frac{2\lambda_{\min}(S^{*})}{\|S^{*}A^{*}-A^{*}S^{*}\|_{2}+2(\|S^{*}\|_{2}+\|A^{*}\|_{2})\,\delta^{(\ell)}}$
such that, for every such $\tau$, the symmetric part of $G_{k}^{(\ell)}$ satisfies $\operatorname{sym}\bigl(G_{k}^{(\ell)}\bigr)\succeq c(\ell)\,I\quad\text{for all }k,$ for some constant $c(\ell)>0$ depending on $\ell$ and the spectral data of $H^{*}$. Moreover, from Lemma 5.1 and the definition of $\alpha_{k}^{(\ell)}$ we also have a uniform bound
$\|G_{k}^{(\ell)}\|_{2}\leq\bigl\|I-\tau\,\alpha_{k}^{(\ell)}\bigr\|_{2}\,\|H^{*}\|_{2}\leq(1+\tau C_{\alpha}(\ell))\|H^{*}\|_{2}=:G_{\max}(\ell)\qquad\text{for all }k,$
for a suitable constant $G_{\max}(\ell)>0$.
For any $z\in\mathbb{R}^{m+n}$,
$\|(I-\eta G_{k}^{(\ell)})z\|_{2}^{2}$ $=\|z\|_{2}^{2}-2\eta\,\bigl\langle\operatorname{sym}(G_{k}^{(\ell)})z,z\bigr\rangle+\eta^{2}\|G_{k}^{(\ell)}z\|_{2}^{2}$
$\leq\bigl(1-2\eta c(\ell)+\eta^{2}G_{\max}(\ell)^{2}\bigr)\,\|z\|_{2}^{2}.$
Choosing
$0<\eta<\bar{\eta}(\ell):=\frac{2c(\ell)}{G_{\max}(\ell)^{2}},$
we obtain $1-2\eta c(\ell)+\eta^{2}G_{\max}(\ell)^{2}<1$, and therefore a uniform bound
$\|I-\eta G_{k}^{(\ell)}\|_{2}\leq q_{0}(\ell)<1\qquad\text{for all }k,$
for some $q_{0}(\ell)\in(0,1)$ depending on $\ell,\tau,H^{*}$ and the constants above.
Since $f,g\in C^{3}(\Omega)$, the game gradient $F$ is of class $C^{2}$. Thus its Jacobian $DF(w)$ is continuous, and in a sufficiently small ball $B_{r_{0}}(w^{*})\subset\Omega$ we can use the Taylor expansion
$F(w)=F(w^{*})+H^{*}(w-w^{*})+R(w),$
with remainder satisfying
$\|R(w)\|_{2}\leq C_{R}\|w-w^{*}\|_{2}^{2},\qquad w\in B_{r_{0}}(w^{*}),$
for some constant $C_{R}>0$.
For $w_{k}\in B_{r_{0}}(w^{*})$ we have $F(w^{*})=0$ and
$w_{k+1}-w^{*}=(w_{k}-w^{*})-\eta\bigl(I-\tau\alpha_{k}^{(\ell)}\bigr)F(w_{k})=(I-\eta G_{k}^{(\ell)})(w_{k}-w^{*})-\eta\bigl(I-\tau\alpha_{k}^{(\ell)}\bigr)R(w_{k}).$
Taking norms and using the bounds above yields
$\|w_{k+1}-w^{*}\|_{2}\leq q_{0}(\ell)\,\|w_{k}-w^{*}\|_{2}+\eta\,\bigl\|I-\tau\alpha_{k}^{(\ell)}\bigr\|_{2}\,\|R(w_{k})\|_{2}.$
By Lemma 5.1 and Corollary 5.2 there exists $C_{I}(\ell)>0$ such that $\|I-\tau\alpha_{k}^{(\ell)}\|_{2}\leq(1+\tau C_{\alpha}(\ell))=:C_{I}(\ell)$ for all $k$, hence
$\|w_{k+1}-w^{*}\|_{2}\leq q_{0}(\ell)\,\|w_{k}-w^{*}\|_{2}+\eta C_{I}(\ell)C_{R}\,\|w_{k}-w^{*}\|_{2}^{2}.$
Choose $0<r(\ell)\leq r_{0}$ such that
$\eta C_{I}(\ell)C_{R}\,r(\ell)\leq\frac{1-q_{0}(\ell)}{2}.$
If $\|w_{k}-w^{*}\|_{2}\leq r(\ell)$, the previous inequality gives
$\|w_{k+1}-w^{*}\|_{2}\leq\Bigl(q_{0}(\ell)+\eta C_{I}(\ell)C_{R}r(\ell)\Bigr)\|w_{k}-w^{*}\|_{2}\leq q(\ell)\,\|w_{k}-w^{*}\|_{2},$
where $q(\ell):=\bigl(1+q_{0}(\ell)\bigr)/2\in(0,1)$. Starting from $w_{0}\in B_{r(\ell)}(w^{*})$ and arguing by induction we obtain that $w_{k}\in B_{r(\ell)}(w^{*})$ for all $k$ and $\|w_{k}-w^{*}\|_{2}\leq q(\ell)^{k}\|w_{0}-w^{*}\|_{2}$, which proves both (1) and (2). ∎
The theorem shows that, for any fixed history length $\ell$, LM–LRSGA inherits the local linear convergence guarantees of LRSGA, up to a change in the admissible stepsizes $\tau$ and $\eta$. The truncation of the memory is reflected only in the constant $\delta^{(\ell)}$ in Corollary 5.2 and hence in the admissible range for $\tau$.
@@ 6.1 Experiment on Mnist
The first experiment is conducted on the MNIST dataset, a classic benchmark in machine learning and computer vision, consisting of 70,000 grayscale images of handwritten digits from 0 to 9. Each image is $28\times 28$ pixels and centered within the frame to standardize positioning. The data set is divided into 60,000 training samples and 10,000 test samples, and its simplicity, small size, and well-defined labels have made it a fundamental starting point for testing and comparing classification algorithms. We used only the training split to train the GAN.
@@ 6.1.1 Results
Figure 1(a) presents the Fréchet Inception Distance (FID) scores over training epochs on the MNIST dataset. It compares our proposed optimizer LM-LRSGA, and its exponential moving average variant (LRSGAEma) against the widely used Adam optimizer. Both LM-LRSGA and LM-LRSGAEma achieve substantially lower FID scores throughout training than Adam, demonstrating superior sample quality. While Adam begins with a significantly higher FID score (above 70) and gradually decreases to around 19 after 100 epochs, LM-LRSGAand LM-LRSGAEma start with lower scores ($\sim 40$) and converge much faster. Notably, LM-LRSGAEma consistently outperforms LM-LRSGA, yielding smoother convergence and the lowest final FID score of approximately 13. These results highlight the effectiveness of our optimizer in stabilizing generative model training and producing higher-quality samples compared to the conventional Adam baseline.
Figure 1(b) shows the generator and discriminator loss trajectories on the MNIST dataset when trained with LM-LRSGA, LM-LRSGAEma, and Adam. The results indicate that LM-LRSGA and LM-LRSGAEma lead to more stable and balanced adversarial training than Adam. With Adam, the generator loss remains consistently lower ($\sim 1.3$) while the discriminator stabilizes around $\sim 0.8$, reflecting a tendency toward imbalance and potential mode collapse. In contrast, LM-LRSGA and LM-LRSGAEma exhibit more synchronized dynamics: the generator loss converges to $\sim 1.4$, while the discriminator loss steadily decreases toward $\sim 0.5$, suggesting healthier competition between the two networks. Notably, LM-LRSGAEma achieves the smoothest convergence, stabilizing both losses earlier and exhibiting minimal oscillation. These findings align with the FID results, demonstrating that our proposed optimizer fosters more stable GAN training and enhances generative performance relative to the Adam baseline.
@@
Figure 1: Comparison of LM-LRSGA, LM-LRSGAEma and Adam optimizers on MNIST datasets.
@@ 6.1.2 Sensitivity analysis
In this section, we analyze the role of the main parameters of the LM-LRSGAEma algorithm. We studied how the model’s ability to generate realistic images changes when we vary the value of a parameter.
Figure 2(a) presents the hyperparameter sensitivity analysis $\eta$ on the minimum Fréchet Inception Distance (FID) score achieved during the experiments. The bar chart shows results for $\eta$ values ranging from 0.05 to 0.3. Performance varies notably across settings, with the highest (worst) FID score observed at $\eta=0.05$, indicating degraded generative quality in this regime. The best results are obtained for $\eta$ values between 0.15 and 0.2, yielding minimum FID scores around 15, suggesting an optimal trade-off between learning stability and image fidelity. Larger $\eta$ values beyond 0.25 slightly worsen the FID, though still outperforming the smallest $\eta$ tested. This pattern highlights the importance of tuning $\eta$ for maximizing GAN performance.
Figure 2(b) reports the hyperparameter sensitivity analysis for the parameter $\tau$. The results are shown for $\tau$ values from 0 to 0.2. The 0 value is an experiment in which the component is ignored. Across this range, the minimum FID scores vary only slightly except for value 0.2, indicating that the model performance is relatively insensitive to $\tau$ within the tested limits. The best generative quality is observed at $\tau=0.002$, both achieving FID scores just above 16, while the highest scores (worst performance) occur around $\tau=0.2$. In general, the flat trend for values below the 0.02 suggests that $\tau$ has a limited effect on GAN performance in this setup only when it is weighted below a threshold.
Figures 2(c) and 2(d) show the results of the experiments to evaluate the effect of $\varepsilon$ and max_history parameters. The charts show only negligible differences between the settings. For the $\varepsilon$ parameter, we expect its impact to be minimal, as it is effectively offset by the residuals observed between consecutive optimization iterations. As for the max_history parameter, we consider a value of 10 sufficient for this dataset to accurately estimate the correction to be applied and the resource usage required to execute the experiment.
@@ 6.2 Experiment on Fashion Mnist
In this section we analyze the results about the Fashion-MNIST dataset, it is a widely used benchmark for image classification tasks, designed as a more challenging drop-in replacement for the original MNIST handwritten digits. It consists of 70,000 grayscale images of size $28\times 28$ pixels, split into 60,000 training samples and 10,000 test samples. We used only the training split to train the GAN models. Each image depicts a single fashion item from one of 10 categories, including T-shirts, trousers, pullovers, dresses, etc. The dataset’s compact size and diversity make it a standard testbed for evaluating machine learning algorithms in vision research.
Figure 3(a) shows the evolution of the FID scores over 100 training epochs for the proposed LM-LRSGA and LM-LRSGAEma optimizer and the Adam baseline optimizer on the Fashion dataset. The FID score decreases sharply for all methods during the initial epochs, dropping from 90 to approximately 55–60 by epoch 20. LM-LRSGA and LM-LRSGAEma consistently achieves lower FID scores for all epochs. Adam achieves comparable performance on the last epoch. We have limited the analysis to 100 epochs to maintain coherence with Mnist experiments.
Figure 3(b) shows the training losses of Adam, LM-LRSGA and LM-LRSGAEma. The same behavior of Mnist dataset is observed in this experiment. The two losses converge to a stable value around the epoch $\sim 30$ for LM-LRSGA and LM-LRSGAEma, while Adam converges to some epochs before. The distance between the two losses confirms the ability of LM-LRSGA and LM-LRSGAEma to better deal than Adam.
@@
Figure 3: Comparison of LM-LRSGA, LM-LRSGAEma, and Adam optimizers on Fashion datasets.